# Search for Stochastic GW Signal as a Complementary Approach to Multi-Higgs Productions at the Hadron Colliders to Probe Dimension Six Operator

### v1

@@ Abstract
We have considered an effective field theory framework in which the Standard Model is extended by a non-renormalizable dimension six operator ($\frac{1}{\Lambda^{2}}(H^{\dagger}H)^{3}$) respecting the symmetries of the Standard Model. Such an operator can affect the dynamics of Higgs field and the electroweak phase transition. Presence of such term modifies the triple Higgs coupling which is already under the lens from di-Higgs searches in various channels at the LHC. Constraint on the scale parameter $\Lambda$ is obtained from the di-Higgs data at the LHC Run II. Additionally, such dimension six terms will give rise to additional couplings which will affect tri-Higgs and four-Higgs productions at future high energy, high luminosity hadron colliders. The shape of the scalar potential also gets modified due to such higher dimensional term composed of Higgs field. It is worthwhile to study the impact of such higher-dimensional operators at suitable gravitational wave (GW) interferometry experiments since GW astronomy has become a good probe for new physics searches. Bounds obtained on the scale parameter $\Lambda$ from the allowed first-order electroweak phase transition is $480\lesssim\Lambda\lesssim 590$ GeV. On the other hand, di-Higgs results from the LHC Run II put a lower bound on $\Lambda\gtrsim 340$ GeV.
@@ 1 Introduction
Meanwhile, the possibility of GW detection in various interferometer based experiments could throw up a nice alternative to test such new physics effect hidden under the dimension six operator under consideration. The effect on the GW spectrum due to $(H^{\dagger}H)^{3}$ term and its possible observation in various GW detectors like, LISA Caprini:2015zlo ; Athron:2023xlk , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi , DECIGO Seto:2001qf ; Kawamura:2006up ; Isoyama:2018rjb have been studied earlier Grojean:2004xa ; Delaunay:2007wb ; Ellis:2018mja  [^1] During the final stage this work, an article Qin:2024dfp with partially similar analysis appeared in the arXiv.. Therefore, time is ripe for new search strategies where signatures from early universe provide substitute for searches in colliders, which are yet to come Profumo:2007wc ; Profumo:2014opa ; Huang:2017jws ; Chen:2017qcz ; Carena:2018vpt ; Alves:2019igs ; Ghosh:2022fzp ; Alves:2020bpi . In addition, one must realize that presence of $(H^{\dagger}H)^{3}$ term in the Lagrangian would, in principle, allows a vertex involving maximum of six Higgses which would contribute to the production of five Higgses in addition to the SM contribution. In this work, we have considered four-Higgs production comprising the five Higgs vertex exclusively arising from SM+H6 scenario.
@@ 2.1 Tree level Potential
where $\mu$ is the bare mass parameter, $\lambda$ is the quartic coupling constant and $\Lambda$ is the scale of new physics, signifying the validity of the theory. From the potential given in Eq.1, it is obvious that $\mu,\Lambda$ each has mass dimension one, whereas, the quartic coupling $\lambda$ is a dimensionless quantity. Considering $H=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ h+v\end{pmatrix}$, one may expand the terms in the potential 1 and estimate required couplings in the Higgs sector. We fix $m_{h}=125$ GeV and $v=246$ GeV, only to consider $\Lambda$ as the free parameter in the theory.
@@ 2.3 Contribution from Ring Diagrams
Figure 1: The central black line in left panel shows the di-Higgs production cross-section measured by the ATLAS collaboration Brigljevic:2024vuv as a function of Higgs self-coupling modifier $\kappa_{\lambda}=\frac{\lambda^{SM+H6}_{hhh}}{\lambda^{SM}_{hhh}}$. The yellow band denotes ±$\pm 2\sigma$ uncertainty around the observed value. The blue curve displays the di-higgs cross-section in SM+H6 scenario. The figure in the right panel shows the allowed region where the theoretical blue curve falls below the experimental $2\sigma$ band. It can be seen that the line goes asymptotically to arbitrary large $\Lambda$ which corresponds to the SM.
@@ 3 Constraints from di-Higgs searches at the LHC Run II
The interferometer-based gravitational wave astronomy experiments is a way to “observe” the stochastic GW relics created during the early universe phase transition. In a very similar manner, one may think of the LHC as a man-made machine to recreate an early universe environment on the earth. Therefore, to measure the properties of the scalar particle discovered at the LHC is of paramount importance in the post-discovery era of the SM-like Higgs boson as it unveils the possible structure of the scalar potential. The Higgs couplings to the SM electroweak gauge bosons are in excellent agreement with the SM expectations, as already confirmed by the LHC Run I and Run II data ATLAS:2023tnc ; ATLAS:2023pwa ; CMS:2022uhn ; CMS:2023gjz . The rarity of events due to extreme smallness of the Higgs couplings to the first two generations of quarks and leptons, the discovery of those couplings can not be established yet at the LHC ATLAS:2022ers ; CMS:2022fxs . However, a circumspect progress has already been made in the couplings to the third generation quarks and lepton ATLAS:2024wfv ; ATLAS:2024moy ; ATLAS:2021qou ; CMS:2022dbt ; CMS:2024fdo . The only other measurable coupling which is also of utmost importance in describing the shape of the scalar potential of the theory is the Higgs self-coupling ($\lambda_{hhh}$). Within the framework of the SM, the value of this coupling is not small at all, however, this self-coupling can be measured only from events where a single Higgs Boson decays into two Higgs Bosons. Additionally, the two Feynman diagrams for the di-Higgs process, the box diagram and the triangle diagram, interfere destructively, resulting in low cross-section for the process. The non-resonant Higgs pair production cross-section for a Higgs mass of $m_{h}=125$ GeV for gluon-gluon-fusion (ggF) process in the SM at next-to-next-to-leading order (NNLO) is $\sigma^{\text{SM}}_{\text{ggF}}=31.1$ fb at the center-of-mass energy of $13$ TeV at the LHC Dawson:1998py ; Shao:2013bz ; deFlorian:2013jea ; deFlorian:2015moa ; Borowka:2016ehy ; Grazzini:2018bsd ; Baglio:2018lrj ; Baglio:2020wgt . The same for vector-boson-fusion (VBF) process is $\sigma^{\text{SM}}_{\text{VBF}}=1.73\pm 0.04$ fb at N3LO in QCD Baglio:2012np ; Frederix:2014hta ; Ling:2014sne ; Dreyer:2018qbw ; Dreyer:2018rfu . In this article, we impose the constraint considering the results from the ATLAS di-Higgs searches ATLAS:2024pov ; ATLAS:2024lsk ; ATLAS:2024ish .
@@ 4 Thermal History of the Universe: SFOPT and GW
For the EWPT to be a SFOPT, the Higgs potential at finite temperature must have a significant barrier between the symmetric phase (where the Higgs field VEV is zero) and the broken phase (where the Higgs field VEV is non-zero). This barrier allows for the nucleation of bubbles of the broken phase within the symmetric phase, which is critical for a strong first-order phase transition. Strong first-order phase transition to occur in the framework of SM, the Higgs boson mass would need to be below about $\sim 70$ GeV Bochkarev:1987wf ; Morrissey:2012db , which is well below the observed mass. However, the dimension six operator $(H^{\dagger}H)^{3}$ augmenting the scalar potential helps in creating the barrier with the observed Higgs mass at $125$ GeV.
@@ 6 Results from Gravitational Wave Astronomy
The GWs can be produced by the collision of the bubbles at a temperature marginally lower than the nucleation temperature, often equated with the $T_{n}$. The parameters quantifying the quality of the GW spectra are:
@@
A large $\alpha$ and a smaller $\beta$ would typically results in good SNR. From Fig.6, we can see that the larger value of the scale parameter $\Lambda$ provides good SNR. Similarly, that a large nucleation temperature is preferred by larger $\Lambda$ can be observed from Fig.7. Therefore, from the Figs. 6 and 7, it can be safely stated that new physics scale on the larger side would give large $\alpha$ and small $\beta/H$ which, in turn, are required for a decent SNR signifying a detectable GW signal at the LISA Caprini:2015zlo ; Athron:2023xlk , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi and DECIGO Seto:2001qf ; Kawamura:2006up ; Isoyama:2018rjb .
Figure 7: The variation of (i) SNR (left panel) and (ii) $\alpha$ (right panel) against $\Lambda$ (in GeV), the parameter of our theory, with the color gradient showing the nucleation temperature $T_{N}$. Large nucleation temperature is preferred by large $\Lambda$. The range is specified by the condition satisfying the SFOPT.
@@
Gravitational waves generated during early universe processes, such as phase transition, typically produce a broad spectrum that can extend over a wide range of frequency. The exact shape of $\Omega_{GW}(f)$ depends on the detailed dynamics of various processes, including bubble collisions, sound waves, and turbulence from phase transitions. For many sources of GWs, the energy density spectrum can be modeled as a power-law function of frequency: $\Omega_{GW}(f)\propto f^{n}$, $n$ being the power-law index. These typically produce a broad spectrum with a peak frequency determined by the temperature and dynamics of the phase transition. The energy density may increase at low frequencies, reach a peak, and then decay at higher frequencies, as can be seen in Fig.8. GW energy density spectrum against frequency for model parameter $\Lambda=580$ GeV is displayed in Fig.8. The individual contributions for turbulence, bubble collision and sound waves are also shown the same plot. The sound wave contribution can be seen to be the dominant one. The peak frequency of the cumulative contribution is well within the reach of LISA Caprini:2015zlo ; Athron:2023xlk , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi and DECIGO Seto:2001qf ; Kawamura:2006up ; Isoyama:2018rjb .
@@ 7 Summary and Conclusions
Figure 8: The GW signal as a function of frequency for $\Lambda=580$ GeV for SM+H6 scenario. The sound wave, collision and turbulence contributions adding to the total energy density spectrum are also shown. The nucleation temperature is around $60$K. It can be seen that the peak frequency is well under the LISA Caprini:2015zlo ; Athron:2023xlk , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi and DECIGO Seto:2001qf ; Kawamura:2006up ; Isoyama:2018rjb detector sensitivities.
The inclusion of the dimension six operator $((H^{\dagger}H)^{3})$ in the Higgs potential provides a concrete scenario where modifications to the Higgs sector can lead to observable consequences in both gravitational wave astronomy and collider experiments. By studying the impact of this operator on the electroweak phase transition and exploring multi-Higgs signatures at colliders, we can gain deeper insights into the nature of the Higgs field and the potential existence of new physics beyond the Standard Model. With no new particles coming up at the LHC, the primary focus of the experimental particle physics has bifurcated into exploring the sky for garnering astrophysical and cosmological data, and various low energy experiments. Furthermore, absence of new resonances has naturally pushed us into thinking in terms of effective theory. The inclusion of dimension six term in the Lagrangian offers some insight of the unknown new physics beyond some energy scale $\Lambda$ at the expense of renormalizability of the theory.
In SM+H6 scenario the triple Higgs and quartic Higgs couplings get modified with respect to the SM. Moreover, it also predicts the five-Higgs and six-Higgs couplings which are the new predictions of this model. Di-Higgs searches at the LHC gives first taste of measurement of the triple Higgs coupling.
@@
We have followed the most well-known formalism of estimating the effective potential by means of loop corrected potentials at zero and finite temperature. On top of that, the requirement of SFOPT to generate detectable stochastic GW from the collision of bubbles of broken EW phase, restricts the scale parameter within $480\leavevmode\nobreak\ \text{GeV}<\Lambda<590\leavevmode\nobreak\ \text{GeV}$.
We have found the numerical values of $\alpha$ and $\beta/H$ for various $\Lambda$ and also computed what is called a good signal in terms of a quantity called SNR (see Fig. 6 and 7 respectively). The peak frequency of resulting GW signal including all three types of contributions can be found (see Fig. 8) to fall within sensitivity reach of LISA Caprini:2015zlo ; Athron:2023xlk , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi and DECIGO Seto:2001qf ; Kawamura:2006up ; Isoyama:2018rjb experiments.

### v2

@@ Abstract
We have considered an effective field theory framework in which the Standard Model is extended by a non-renormalizable dimension six operator ($\frac{1}{\Lambda^{2}}(H^{\dagger}H)^{3}$) respecting the symmetries of the Standard Model. Such an operator can affect the dynamics of Higgs field and the electroweak phase transition. Presence of such term modifies the triple Higgs coupling which is already under the lens from di-Higgs searches in various channels at the LHC. Constraint on the scale parameter $\Lambda$ is obtained from the di-Higgs data at the LHC Run II. Additionally, such dimension six terms will give rise to additional couplings which will affect tri-Higgs and four-Higgs productions at future high energy, high luminosity hadron colliders. The shape of the scalar potential also gets modified due to such higher dimensional term composed of Higgs field. It is worthwhile to study the impact of such higher-dimensional operators at suitable gravitational wave (GW) interferometry experiments since GW astronomy has become a good probe for new physics searches. Bounds obtained on the scale parameter $\Lambda$ from the allowed first-order electroweak phase transition is $450\leavevmode\nobreak\ \text{GeV}\lesssim\Lambda\lesssim 590$ GeV. On the other hand, di-Higgs results from the LHC Run II put a lower bound on $\Lambda\gtrsim 340$ GeV.
@@ 1 Introduction
Meanwhile, the possibility of GW detection in various interferometer based experiments could throw up a nice alternative to test such new physics effect hidden under the dimension six operator under consideration. The effect on the GW spectrum due to $(H^{\dagger}H)^{3}$ term and its possible observation in various GW detectors like, LISA Caprini:2015zlo ; Athron:2023xlk , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi , DECIGO Seto:2001qf ; Kawamura:2006up ; Isoyama:2018rjb have been studied earlier Grojean:2004xa ; Delaunay:2007wb ; Ellis:2018mja ; Banerjee:2024qiu  [^1] During the final stage this work, an article Qin:2024dfp with partially similar analysis appeared in the arXiv.. Therefore, time is ripe for new search strategies where signatures from early universe provide substitute for searches in colliders, which are yet to come Profumo:2007wc ; Profumo:2014opa ; Huang:2017jws ; Chen:2017qcz ; Carena:2018vpt ; Alves:2019igs ; Ghosh:2022fzp ; Alves:2020bpi ; Roy:2022gop . In addition, one must realize that presence of $(H^{\dagger}H)^{3}$ term in the Lagrangian would, in principle, allows a vertex involving maximum of six Higgses which would contribute to the production of five Higgses in addition to the SM contribution. In this work, we have considered four-Higgs production comprising the five Higgs vertex exclusively arising from SM+H6 scenario.
@@ 2.1 Tree level Potential
where $\mu$ is the bare mass parameter, $\lambda$ is the quartic coupling constant and $\Lambda$ is the scale of new physics, signifying the validity of the theory. From the potential given in Eq.1, it is obvious that $\mu,\Lambda$ each has mass dimension one, whereas, the quartic coupling $\lambda$ is a dimensionless quantity. Considering $H=\frac{1}{\sqrt{2}}\begin{pmatrix}0\\ h+v\end{pmatrix}$, one may expand the terms in the potential 1 and estimate required couplings in the Higgs sector. From the three parameters in eqn.1, we trade in $\mu^{2}$ and $\lambda$ using the minimization condition and mass determining condition, namely, $V^{{}^{\prime}}_{0}(h=v)=0$ and $V^{{}^{\prime\prime}}_{0}(h=v)=m^{2}_{h}$ to obtain the following expressions Delaunay:2007wb ; Ellis:2018mja :
$\mu^{2}=\frac{m^{2}_{h}}{2}-\frac{3v^{4}}{4\Lambda^{2}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \lambda=\frac{m^{2}_{h}}{2v^{2}}-\frac{3v^{2}}{2\Lambda^{2}}$
Therefore, we fix $m_{h}=125$ GeV and $v=246$ GeV, only to consider $\Lambda$ as the free parameter in the theory.
@@ 2.3 Contribution from Ring Diagrams
Figure 1: The central black line in left panel shows the di-Higgs production cross-section measured by the ATLAS collaboration ATLAS:2021tyg as a function of Higgs self-coupling modifier $\kappa_{\lambda}=\frac{\lambda^{SM+H6}_{hhh}}{\lambda^{SM}_{hhh}}$. The yellow band denotes ±$\pm 2\sigma$ uncertainty around the observed value. The blue curve displays the di-higgs cross-section in SM+H6 scenario. The figure in the right panel shows the allowed region where the theoretical blue curve falls below the experimental $2\sigma$ band. It can be seen that the line goes asymptotically to arbitrary large $\Lambda$ which corresponds to the SM.
@@ 3 Constraints from di-Higgs searches at the LHC Run II
The interferometer-based gravitational wave astronomy experiments is a way to “observe” the stochastic GW relics created during the early universe phase transition. In a very similar manner, one may think of the LHC as a man-made machine to recreate an early universe environment on the earth. Therefore, to measure the properties of the scalar particle discovered at the LHC is of paramount importance in the post-discovery era of the SM-like Higgs boson as it unveils the possible structure of the scalar potential Agrawal:2019bpm . The Higgs couplings to the SM electroweak gauge bosons are in excellent agreement with the SM expectations, as already confirmed by the LHC Run I and Run II data ATLAS:2023tnc ; ATLAS:2023pwa ; CMS:2022uhn ; CMS:2023gjz . The rarity of events due to extreme smallness of the Higgs couplings to the first two generations of quarks and leptons, the discovery of those couplings can not be established yet at the LHC ATLAS:2022ers ; CMS:2022fxs . However, a circumspect progress has already been made in the couplings to the third generation quarks and lepton ATLAS:2024wfv ; ATLAS:2024moy ; ATLAS:2021qou ; CMS:2022dbt ; CMS:2024fdo . The only other measurable coupling which is also of utmost importance in describing the shape of the scalar potential of the theory is the Higgs self-coupling ($\lambda_{hhh}$). Within the framework of the SM, the value of this coupling is not small at all, however, this self-coupling can be measured only from events where a single Higgs Boson decays into two Higgs Bosons. Additionally, the two Feynman diagrams for the di-Higgs process, the box diagram and the triangle diagram, interfere destructively, resulting in low cross-section for the process. The non-resonant Higgs pair production cross-section for a Higgs mass of $m_{h}=125$ GeV for gluon-gluon-fusion (ggF) process in the SM at next-to-next-to-leading order (NNLO) is $\sigma^{\text{SM}}_{\text{ggF}}=31.1$ fb at the center-of-mass energy of $13$ TeV at the LHC Dawson:1998py ; Shao:2013bz ; deFlorian:2013jea ; deFlorian:2015moa ; Borowka:2016ehy ; Grazzini:2018bsd ; Baglio:2018lrj ; Baglio:2020wgt . The same for vector-boson-fusion (VBF) process is $\sigma^{\text{SM}}_{\text{VBF}}=1.73\pm 0.04$ fb at N3LO in QCD Baglio:2012np ; Frederix:2014hta ; Ling:2014sne ; Dreyer:2018qbw ; Dreyer:2018rfu . In this article, we impose the constraint considering the results from the ATLAS di-Higgs searches ATLAS:2024pov ; ATLAS:2024lsk ; ATLAS:2024ish .
@@ 4 Thermal History of the Universe: SFOPT and GW
For the EWPT to be a SFOPT, the Higgs potential at finite temperature must have a significant barrier between the symmetric phase (where the Higgs field VEV is zero) and the broken phase (where the Higgs field VEV is non-zero). This barrier allows for the nucleation of bubbles of the broken phase within the symmetric phase, which is critical for a strong first-order phase transition. Strong first-order phase transition to occur in the framework of SM, the Higgs boson mass would need to be below about $\sim 70$ GeV Bochkarev:1987wf ; Morrissey:2012db , which is well below the observed mass. However, the dimension six operator $(H^{\dagger}H)^{3}$ augmenting the scalar potential helps in creating the barrier with the observed Higgs mass at $125$ GeV. Note that only a small range of $\Lambda$ is favourable for EWPT as can be observed from tab.1. For $590450\leavevmode\nobreak\ \text{GeV}<\Lambda<$ 590 GeV, EWPT proceeds via successful nucleation of bubbles of broken phase of electroweak symmetry. $\Lambda$ beyond these ranges are not conducive for EWPT.
$\Lambda$ (GeV) $\alpha$ $\beta/H_{n}$ $T_{n}$ (GeV)
590 0.2909 215.38 39.14
550 0.0528 1608.20 97.32
500 0.0232 3922.72 147.10
450 0.0069 9738.97 204.24
Table 1: Variations of two dimensionless parameters $\alpha$ and $\beta/H_{n}$, quantifying the quality of GW spectra (defined in sec. 6), and the nucleation temperature $T_{n}$ against the energy scale $\Lambda$. For $\Lambda\lesssim 450$ GeV and $\Lambda\gtrsim 590$ GeV, bubble nucleation does not take place which means EWPT does not occur beyond these values of $\Lambda$s.
@@ 6 Results from Gravitational Wave Astronomy
The GWs can be produced by the collision of the bubbles at a temperature marginally lower than the nucleation temperature, often equated with the $T_{n}$. The (dimensionless) parameters quantifying the quality of the GW spectra are:
@@
A large $\alpha$ and a smaller $\beta$ would typically results in good SNR. From Fig.6, we can see that the larger value of the scale parameter $\Lambda$ provides good SNR. However, that a smaller nucleation temperature is preferred by larger $\Lambda$ can be observed from Fig.7. Therefore, from the Figs. 6 and 7, it can be safely stated that new physics scale on the larger side would give large $\alpha$ and small $\beta/H$ which, in turn, are required for a decent SNR signifying a detectable GW signal at the LISA Caprini:2015zlo ; Athron:2023xlk , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi and DECIGO Seto:2001qf ; Kawamura:2006up ; Isoyama:2018rjb .
Figure 7: The variation of (i) SNR (left panel) (at LISA) and (ii) $\alpha$ (right panel) against $\Lambda$ (in GeV), the parameter of our theory, with the color gradient showing the nucleation temperature $T_{N}$. Large nucleation temperature is preferred by large $\Lambda$. The range is specified by the condition satisfying the SFOPT.
@@
Figure 8: The GW signal as a function of frequency for $\Lambda=580$ GeV for SM+H6 scenario. The sound wave, collision and turbulence contributions adding to the total energy density spectrum are also shown. The nucleation temperature is around $60$ GeV. It can be seen that the peak frequency is well under the LISA Caprini:2015zlo ; Athron:2023xlk , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi and DECIGO Seto:2001qf ; Kawamura:2006up ; Isoyama:2018rjb detector sensitivities.
Gravitational waves generated during early universe processes, such as phase transition, typically produce a broad spectrum that can extend over a wide range of frequency. The exact shape of $\Omega_{GW}(f)$ depends on the detailed dynamics of various processes, including bubble collisions, sound waves, and turbulence from phase transitions. For many sources of GWs, the energy density spectrum can be modeled as a power-law function of frequency: $\Omega_{GW}(f)\propto f^{n}$, $n$ being the power-law index. These typically produce a broad spectrum with a peak frequency determined by the temperature and dynamics of the phase transition. The energy density may increase at low frequencies, reach a peak, and then decay at higher frequencies, as can be seen in Fig.8. GW energy density spectrum against frequency for model parameter $\Lambda=580$ GeV is displayed in Fig.8. The individual contributions for turbulence, bubble collision and sound waves are also shown the same plot. The sound wave contribution can be seen to be the dominant one. The peak frequency of the cumulative contribution is well within the reach of LISA Caprini:2015zlo ; Athron:2023xlk , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi and DECIGO Seto:2001qf ; Kawamura:2006up ; Isoyama:2018rjb . In Fig.9, we have displayed the shift in the peak frequency of GW spectra for different benchmark values of the scale parameter $\Lambda$ for SM+H6 scenario. We observe larger $\Lambda$ favours lower nucleation temperature ($T_{n}$) (and equivalently, smaller peak frequency) which can also be corroborated by the Fig.7. The nucleation temperatures $T_{n}=137.2,97.3,60.2,39.1$ (all in GeV) used in Fig.9 correspond to $\Lambda=510,550,580,590$ (all in GeV).
@@ 7 Summary and Conclusions
Figure 9: The GW spectra as a function of frequency for some benchmark values of the scale parameter $\Lambda$ for SM+H6 scenario. The curves corresponding to lower nucleation temperatures ($T_{n}$) or, equivalently lower peak frequencies correspond to higher values of $\Lambda$. The integrated detector sensitivities of LISA Caprini:2015zlo ; Athron:2023xlk , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi and DECIGO Seto:2001qf ; Kawamura:2006up ; Isoyama:2018rjb are also shown. All $T_{n}$s in the figure are in GeV.
The absence of new resonances has naturally pushed us into thinking in terms of effective theory. The inclusion of a dimension six operator $((H^{\dagger}H)^{3})$ in the Lagrangian offers some insight of the unknown new physics beyond some energy scale $\Lambda$ at the expense of renormalizability of the theory. Such term in the Higgs potential provides a concrete scenario where modifications to the Higgs sector can lead to observable consequences in both gravitational wave astronomy and collider experiments. By studying the impact of this operator on the electroweak phase transition and exploring multi-Higgs signatures at colliders, we can gain deeper insights into the nature of the Higgs field and the potential existence of new physics beyond the Standard Model. In SM+H6 scenario the triple Higgs and quartic Higgs couplings get modified with respect to the SM. Triple Higgs coupling is already under probe through the di-Higgs searches at the LHC. Moreover, SM+H6 model also predicts five-Higgs and six-Higgs couplings which are absent in the SM.
@@
We have followed the most well-known formalism of estimating the effective potential by means of loop corrected potentials at zero and finite temperature. On top of that, the requirement of SFOPT to generate detectable stochastic GW from the collision of bubbles of broken EW phase, restricts the scale parameter within $450\leavevmode\nobreak\ \text{GeV}\lesssim\Lambda\lesssim 590$ GeV.
We have found the numerical values of $\alpha$ and $\beta/H$ for various $\Lambda$ and also computed what is called a good signal in terms of a quantity called SNR (see Fig. 6 and 7 respectively). The peak frequency of resulting GW signal including all three types of contributions can be found (see Fig. 8) to fall within sensitivity reach of LISA Caprini:2015zlo ; Athron:2023xlk , BBO Crowder:2005nr ; Corbin:2005ny ; Harry:2006fi and DECIGO Seto:2001qf ; Kawamura:2006up ; Isoyama:2018rjb experiments. Higher new physics scale favours lower nucleation temperature, and hence, satisfies the condition for strongly supercooled state in the early universe, which is required for stronger GW spectrum (see Fig.9).